Find the equation of a line that passes through (-3,-2), (3,6) using slope intercept form. I got y=4/3x+3 but the book says the answer is y=4/3x+2. How did they get that?
Answer by Alan3354(69443)

Find the equation of a line that passes through (-3,-2), (3,6) using slope intercept form. I got y=4/3x+3 but the book says the answer is y=4/3x+2. How did they get that?
-----------
Slope m = (-2-6)/(-3-3) = 4/3 --- no problem there.
---------------
y-y1 = m*(x-x1) where (x1,y1) is either point
y+2 = (4/3)*(x+3)
y+2 = (4/3)x + 4
y = (4/3)x + 2